OpenStudy (anonymous):

Can anyone solve 5h-9=-16+6h

OpenStudy (tkhunny):

You can. Subtract 5h. Then what?

geerky42 (geerky42):

Both sides, just saying.

OpenStudy (anonymous):

I know its 7 but not sure if its -7 or 7

OpenStudy (anonymous):

Goal: To get "h" by itself on one side of the equation How: Adding / Subtracting / Multiplying / Dividing numbers from BOTH sides of the equation. We do this by adding the opposite and dividing by coefficients.

geerky42 (geerky42):

We need to isolate h, so subtract both sides by 5h, then add both sides by 16. There you go.

geerky42 (geerky42):

Is this clear? @bgenevieve96

OpenStudy (anonymous):

7

geerky42 (geerky42):

Yes!

OpenStudy (anonymous):

Thanks just wanted to double check

geerky42 (geerky42):

I may be wrong, so you may need to triple check. ;)

OpenStudy (anonymous):

Haha thanks

OpenStudy (tkhunny):

Seriously? It's mathematics, folks. Let's quite guessing. 5h-9=-16+6h Subtract 5h 5h-9-5h=-16+6h-5h 5h-5h-9=-16+6h-5h -9 = -16 + h Add 16 -9 + 16 = -16 + h + 16 7 = -16 + 16 + h 7 = h No more guessing. Be clear, consistent, and correct.

geerky42 (geerky42):

"Let's quite guessing." Seriously?

OpenStudy (tkhunny):

Let's also quit spelling badly.